OpenStudy (anonymous):

The perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 9 inches and its preimeter is 44 inches, find the length of the rectangle.

Parth (parthkohli):

 \[9 + 9 + ? +? =44\]

OpenStudy (anonymous):

okay, well 9+9=18 so would I then multiply 18 into 44?

OpenStudy (anonymous):

no divide

Parth (parthkohli):

No, it's the sum of four sides

OpenStudy (anonymous):

oh so add.

OpenStudy (anonymous):

so then how do I figure out the lenght just start plugging in numbers until I get 44?

OpenStudy (unklerhaukus):

(44-18 )/ 2 =

OpenStudy (anonymous):

oh oaky that makes more sense

OpenStudy (anonymous):

so the length is 13

OpenStudy (unklerhaukus):

does 9+9+13+13= 44?

OpenStudy (anonymous):

yes it does.

OpenStudy (anonymous):

Thank you for your help